kuKSSkudkhHf:

What is the slope-intercept form of the equation of the line that passes through the point (–6, 1) and is perpendicular to the graph of 2x + 3y = –5?

xxDeppressionxx:

Do you have any question answer choices?

Mercury:

first find the slope of 2x + 3y = –5 re-write the equation in y = mx + b form by solving for y, and seeing what number m is being multiplied to x from there, since it's asking for a line that's perpendicular, take the opposite reciprocal (a.k.a -1/m). this is the slope of your new line. now, since it's asking for slope intercept form (a.k.a y = mx + b), take the slope you just calculated, along with the point (-6,1) and plug the appropriate quantities into y = mx + b to solve for b finally, take your m and b value and plug it into the generic equation y = mx + b